OpenStudy (anonymous):

Solve for w 9w – 6 = -3(w – 2) Simplify your answer as much as possible

OpenStudy (anonymous):

if anybody solve this right I will give you a medal

OpenStudy (hesan):

9w - 6 = -3w + 6 9w + 3w = 6 + 6 12w = 12 w = 12/12 w=1